# Diffusion-based subsurface multiphysics monitoring and forecasting

### v2

@@ Abstract
  • We develop a novel monitoring and forecasting framework using video diffusion models of subsurface changes due to, for example, CO2 injection.
  • This video generation framework can be a simulation engine for multiphysics evolution, as well as maintain a physical consistency between different physical parameters, thus enabling subsequent forecasting and monitoring.
  • The proposed framework is successfully applied to multiphysics monitoring and forecasting for permeability, CO2 saturation, velocity, density, and reverse time migration (RTM) images.
@@ 1 Introduction
Carbon capture and storage (CCS) play a crucial role in mitigating greenhouse gas emissions, particularly from industrial outputs. This significance is underscored in the recent International Panel on Climate Change report [IPCC, 2022]. CCS involves capturing carbon dioxide (CO2), transporting it, and injecting it into subsurface storage sites. However, to ensure its effectiveness and mitigate associated risks, monitoring, verification, and accounting of the injected CO2 plume are essential [Sun et al., 2023]. Effective monitoring hinges on accurate and robust systems. Seismic monitoring, as highlighted by Lumley [2010], can provide such a system that is both accurate and covers a large area. It operates on the principle that CO2 injection alters the elastic properties of the subsurface, such as velocities and densities. These changes can be inferred, to some extent, from the imaging and inversion of seismic recordings [Bosch et al., 2010, Hicks et al., 2016, Li and Alkhalifah, 2022, Hu et al., 2023]. However, conventional methods like statistical (e.g., Monte Carlo simulation) or wave equation-based approaches are computationally demanding and thus costly for realistic applications. In addition to the monitoring, predicting the future states of CO2 flow is also important as they can provide information on how much CO2 can be stored safely and help us optimize our decisions with respect to injections. Using the above conventional methods, we struggle to track the CO2 plume evolution and predict future subsurface states. Solving non-linear partial differential equations for a multi-phase flow phenomenon [Pruess and Nordbotten, 2011] can help us predict future states based on historical states, but this prediction is still computationally expensive [Witte et al., 2023].
In the era of artificial intelligence, modern machine-learning techniques have opened new avenues to tackle the above issues. Recent studies [Li et al., 2018, Zhong et al., 2019, Zhou et al., 2019, Li et al., 2020, Wen et al., 2021, 2022, Witte et al., 2023, Louboutin et al., 2023b, Yin et al., 2023b, Sun et al., 2023, Stepien et al., 2023, Gahlot et al., 2023] have leveraged machine learning techniques for various aspects of CCS monitoring and forecasting. Yet, these methods primarily focus on leakage detection, accelerating CO2 plume simulation and the CCS modeling, single physics and 2D mapping with generative models, or cascade multiphysics inversion. Real-world monitoring and forecasting of subsurface CO2 injections involve a complex multiphysics process encompassing fluid flow simulation, rock physics, and wave propagation.
@@
To achieve an end-to-end multiphysics inversion and maintain the consistency of the evolution of different variables in terms of physical laws, we introduce a novel diffusion model-based framework for monitoring and forecasting in this paper. Instead of using 2D denoising diffusion probabilistic models (DDPMs) [Ho et al., 2020] to learn the inverse mapping between the seismic gather shots and the CO2 plume shape [Sun et al., 2023], our framework trains a diffusion model to learn the data distribution of 3D multiphysics evolution (spanning 2D spatial and 1D time-lapse domains). This 3D volume incorporates variables such as permeability, CO2 saturation, velocity, density, and reverse time migration (RTM) images. The philosophy behind the use of 3D volume (video) prediction is that the physical process in the real world can be regarded as video sequences. Thus, the forecasting and inversion process will be thought of as the tasks for video generation given the history frames or partial information. Unlike recursively predicting the future states along the time axis using a neural operator, which might have errors accumulation [Ding et al., 2024], the video diffusion model can simultaneously generate multistep future states. During inference, we employ reconstruction guidance to facilitate forecasting and inversion using historical frames or observed RTM images. Thanks to the probabilistic nature of the diffusion model, we can easily perform uncertainty analysis for forecasting and inversion. In this paper, we validate our method using the 3D Compass model [Jones et al., 2012].
@@
  • We propose a monitoring and forecasting framework using video diffusion models.
@@
The remainder of the paper is organized as follows. In section 2, we introduce the video diffusion models and the proposed framework for forecasting and monitoring based on video diffusion models with reconstruction guidance. Then, we share our numerical simulation for the data generation. In section 3, we present numerically test the proposed method on compass models, including the results of unconditional generation, monitoring, and forecasting. Finally, we discuss and summarize the work, and present insights in sections 4 and 5.
@@ 2.2 Video Diffusion models
Denoising diffusion probabilistic models (DDPMs) are a family of generative models that enable the generation of samples from a given distribution. They have shown a state-of-the-art image and video synthesis performance among other generative models, such as variational autoencoders [Kingma and Welling, 2013], generative adversarial networks [Goodfellow et al., 2014], normalizing flows [Rezende and Mohamed, 2015]. The general idea behind DDPMs is to train a neural network to learn to transform a prescribed distribution (often a Gaussian distribution) to an approximation of the target data distribution. The pipeline of the diffusion model includes two processes: training (a.k. forward process from $\mathbf{x}_{0}$ to $\mathbf{x}_{T-1}$) and sampling (a.k. reverse process from $\mathbf{x}_{T-1}$ to $\mathbf{x}_{0}$), shown in Figure 1. The subscript $T$ of the $\mathbf{x}_{T}$ here denotes the time step in the diffusion process. Our target is to fit the data distribution of spatial-temporal evolution for multiple variables, which are 3D volumes. We process it as videos and thus use video diffusion models [Ho et al., 2022].
@@
Figure 1: Workflow of the diffusion process and its reverse process, and the reconstruction guided generation for forecasting and inversion. The yellow box denotes the scaled learned score function.
@@ 2.5 CO2 related Multiphysics datasets generation
Fluid flow simulation. Before we can simulate fluid flow, we need to define the permeability $K$ of the subsurface model. Here, as suggested by Yin et al. [2023b], we convert the wave speed of the compass model to log-permeability values to make up alternating high- and low-permeability layers in the reservoir with a seal on top. The porosity used here is a constant value of 0.25. Then, we place one injection well in the deeper part of the target domain and consider a constant injection rate for the experiment here. After the injection, we use a two-phase simulation to obtain the CO2 saturation model in Figure 2 with softwares Jutul.jl [Møyner et al., 2023] and JutulDarcyRules.jl [Yin et al., 2023a].
Rock physics modeling. Since we use seismic methods to monitor the subsurface CO2 changes, we need to obtain the velocity changes resulting from injection and the changes in saturation. We use the patchy saturation model [Louboutin et al., 2023b] to convert the saturation model to velocity and density models, in which an increase in CO2 saturation will result in a decrease in velocity and density.
Seismic modeling and reverse time imaging. After obtaining the seismic properties changes provided by the velocity and density, we simulated the seismic response based on the impedance using born modeling and applied RTM to obtain images [Yin et al., 2023b] using the software JUDI.jl [Louboutin et al., 2023a]. The details on the source and receiver locations will be shared in the examples section. As the seismic responses to the velocity and density changes are quite small compared to the baseline monitoring, we do not use the RTM images directly. Instead, we use the residuals between the current RTM image and the baseline RTM image as the seismic observation. The Born modeling here will secure for us clean images as if processing and demultiple were applied to conventionally observed data.
@@ 2.6 The backbone network and training configuration
The backbone of the video diffusion model can have any architecture, e.g., 3D-Unet based [Cicek et al., 2016] or transformer based [Peebles and Xie, 2022]. In this paper, since we do not use a variational autoencoder to compress the videos to the latent space, using a transformer would be costly due to the quadratic computational complexity in the attention mechanism of the transformer. Thus, we use a 3D-Unet, which is factorized over space and time. Compared to the 2D Unet often used for image processing, whose convolutional layer is often with the kernel size of 3$\times$3, we use 1$\times$3$\times$3 convolutional kernel for the video. The dimensions of the kernel correspond to the time frames, height, and width of the multiphysics evolution. The first attention block used in the 3D-Unet applies to the space dimension only, in which the time frame axis is regarded as the batch dimension, combined with a following temporal attention block. To distinguish the order of the frames (the time order of the multiphysics evolution), the relative position embeddings [Shaw et al., 2018] are used in each temporal attention block. This paper’s backbone network of the video diffusion model is designed to consider five channels, including permeability, CO2 saturation, velocity, density, and RTM images.
@@ 3 Numerical Experiments
In our experiments, we employed the 3D Compass velocity model, a synthetic model that closely mimics the geological characteristics found in the southeastern region of the North Sea. This model serves as an ideal testbed due to its realistic representation of the subsurface geological features. We extract 800 two-dimensional slices from the three-dimensional compass velocity model, having dimensions of $64\times 64$, with an interval of 25 $m$ in the horizontal direction and 12 $m$ in the vertical direction. In this paper, we do not consider the leakage scenario and assume the permeability is constant over time, though this scenario can also be incorporated if needed. For the fluid flow simulation, we control the CO2 injection rate to achieve a 10% storage capacity over an 18-year period. The constant CO2 injection happens at the fixed well location (often known in field experiments). From this simulation, we sampled six frames starting from the third year with a time-lapse interval of three years. These frames were crucial in analyzing the temporal evolution of the injected CO2. Subsequently, we applied rock physics modeling to calculate the corresponding changes in velocities and densities due to CO2 injection for each CO2 saturation at different times. We collected the simulated seismic data for each velocity and density model by placing six sources on the surface with a 24 Hz Ricker wavelet and 200 receivers on the sea bottom. The recordings lasted for 1 $s$ with a sampling rate of 4 $ms$. Then, we employ RTM to compute the time-lapse seismic image changes by subtracting them from the baseline image. In the video diffusion training phase, we used an Adam optimizer with a learning rate of 1e-4, a batch size of 64, and a maximum of 150,000 iterations until the neural network converged, as evidence by the training loss curve, shown in Figure 3.
@@ 3.2 Forecasting
As mentioned in Section 2, this well-trained unconditional video diffusion model can be used for forecasting and inversion. We first test it for the forecasting task to evaluate the ability to predict future CO2 evolution based on the history frames as that can help us evaluate the capacity for CO2 storage for decision-making optimization. With a reconstruction scale $w_{r}$ of 1.0, the conditional sampling results are shown in Figure 5. Given the observations and records of the subsurface states in years 3, 6, 9, we forecast the future states for years 12, 15, and 18. They are consistent and reasonable in their evolution and generally maintain the coupling relationship between the five physical parameters. To quantitate the forecasting errors for each variable, we compare them with the reference numerical simulations, shown in Figure 6. The errors are small and are focused mainly in the region of the CO2 plume. This generally validates our model’s accuracy and highlights its potential as a powerful tool for strategic planning in CO2 storage projects.
@@
Figure 5: Forecasting generation (b) given three history frames (a), which are denoted by the red box. The rows represent properties described in Figure 4.
@@
Figure 7: Inversion results (b) given three history frames and observed-data-based RTM images (a), which are denoted by the red box. The rows represent properties described in Figure 4

### v3

# Diffusion-based subsurface CO2 multiphysics monitoring and forecasting

@@ Abstract
  • We develop a novel monitoring and forecasting framework using video diffusion models of subsurface changes due to, e.g., CO2 injection.
  • This approach can be a simulation engine for multiphysics evolution and maintain physical consistency between multiple variables.
  • The proposed method is applied to multiphysics monitoring and forecasting for permeability, CO2 saturation, velocity, density, and RTM images.
@@ 1 Introduction
Carbon capture and storage (CCS) play a crucial role in mitigating greenhouse gas emissions, particularly from industrial outputs. This significance is underscored in the recent International Panel on Climate Change report [IPCC, 2022]. CCS involves capturing carbon dioxide (CO2), transporting it, and injecting it into subsurface storage sites. However, to ensure its effectiveness and mitigate associated risks, monitoring, verification, and accounting of the injected CO2 plume are essential [Sun et al., 2023]. Effective monitoring hinges on accurate and robust systems. Seismic monitoring, as highlighted by Lumley [2010], can provide such a system that is both accurate and covers a large area. It operates on the principle that CO2 injection alters the elastic properties of the subsurface, such as velocities and densities. These changes can be inferred, to some extent, from the imaging and inversion of seismic recordings [Bosch et al., 2010, Hicks et al., 2016, Li and Alkhalifah, 2022, Hu et al., 2023]. However, conventional methods like statistical (e.g., Monte Carlo simulation) or wave equation-based approaches are computationally demanding and thus costly for realistic applications. In addition to the monitoring, predicting the future states of CO2 flow is also important as they can provide information on how much CO2 can be stored safely and help us optimize our decisions with respect to injections. Using the above conventional methods, we struggle to track the CO2 plume evolution and predict future subsurface states. Solving non-linear partial differential equations for a multi-phase flow phenomenon [Pruess and Nordbotten, 2011, Ringrose, 2020] can help us predict future states based on historical states. Although accurate and generic, this prediction is still computationally expensive because of the formulations involved in the iterative solvers [Blunt et al., 1992, Graupner et al., 2011], especially for the inversion application involved with a repeated simulation [Witte et al., 2023].
In the era of artificial intelligence, modern machine-learning techniques have opened new avenues, data-driven approaches, to tackle the above issues. Recent studies have leveraged machine learning techniques for various aspects of CCS monitoring and forecasting. Li et al. [2018] proposed the use of an artificial neural network to predict the probability of long-term leakage of wells in CO2 sequestration operations. Zhou et al. [2019] proposed to use a spatial-temporal densely connected convolutional neural networks to detect CO2 leakage from seismic data, where the network design is geared to capture spatial and temporal characteristics of seismic data. Beyond leakage detection, Zhong et al. [2019] proposed a conditional deep convolution generative adversarial network as a surrogate model to predict CO2 plume migration. Stepien et al. [2023] enhanced this approach by predicting CO2 plume migration with observations at three well locations. Wen et al. [2021] proposed the CCSNet to efficiently handle an entire class of multiphase flow problems, providing predictions of CO2 gas saturation distribution, pressure buildups, the molar fractions of CO2 and fluid densities for gas and liquid phases, in CCS modeling. Wen et al. [2022] further accelerate the machine learning-based prediction using Fourier neural operators [Li et al., 2020b]. Witte et al. [2023] applied wavelet neural operators to enable fast CO2 saturation simulation on large-scale geomodels. To incorporate more observation to better monitor the CO2 migration and optimize reservoir management, Li et al. [2020a] proposed a coupled time-lapse full waveform inversion framework to estimate subsurface properties such as rock permeability and porosity from seismic data by intrusive automatic differentiation. Louboutin et al. [2023b] further expanded this cascaded multiphysics inversion with machine learning priors, which represent the permeability using normalizing flow [Rezende and Mohamed, 2015]. Based on this cascaded multiphysics framework, Yin et al. [2023b] developed an explainable leakage detection network from time-lapse seismic data. Beyond cascaded multiphysics inversion, Sun et al. [2023] directly learn the mapping from seismic shot gathers to the CO2 plume using denoising diffusion probabilistic models. Gahlot et al. [2023] proposed to use conditional normalizing flows to obtain the CO2 saturation based on the available seismic images and well log observations. Chen et al. [2024b] proposed to treat the CO2 monitoring task as a segmentation task to save the interpretation time. Yet, these methods primarily focus on leakage detection and CO2 plume interpretation, accelerating CO2 plume simulation and CCS modeling, single physics and 2D mapping with generative models, or a cascaded multiphysics inversion. Real-world monitoring and forecasting of subsurface CO2 injections involve a complex multiphysics process encompassing fluid flow simulation, rock physics, and wave propagation.
@@
To achieve an end-to-end multiphysics inversion and maintain the consistency of the evolution of different variables in terms of physical laws, we introduce a novel generative model-based framework for monitoring and forecasting in this paper. Rather than directly performing the deterministic prediction or the cascade inversion, we train the model to represent the joint distribution among multiple physical variables and perform the monitoring and forecasting with conditional sampling in this distribution. Specifically, we chose denoising diffusion probabilistic models (DDPMs) [Ho et al., 2020] to represent the data distribution. Two-dimensional DDPMs have shown promising results in seismic processing [Durall et al., 2023], diffusion regularized full waveform inversion [Wang et al., 2023], the inverse mapping between the seismic shot gathers and the CO2 plume shape [Sun et al., 2023], controllable velocity synthesis [Wang et al., 2024], and seismic diffraction separation and imaging [Zhang et al., 2024]. Instead of a 2D version, our framework trains a diffusion model to learn the data distribution of 3D multiphysics evolution (spanning 2D along the spatial domain and 1D along time-lapse direction). Those 3D volumes incorporate variables such as permeability, CO2 saturation, velocity, density, and subsurface seismic imaging results using reverse time migration (RTM). The philosophy behind the use of 3D volume (video) prediction is that the physical process in the real world can be regarded as video sequences. Thus, the forecasting and inversion process will be thought of as the tasks for video generation given the history frames or partial information. Unlike recursively predicting the future states along the time axis using a neural operator, which might have errors accumulation [Ding et al., 2024], the video diffusion model can simultaneously generate multistep future states without the need for sequential inference. During inference, we employ reconstruction guidance to facilitate forecasting and inversion using historical frames or observed RTM images. Thanks to the probabilistic nature of the diffusion model, we can easily perform uncertainty analysis for forecasting and inversion. In this paper, we validate our method using the models extracted from the 3D Compass model [Jones et al., 2012].
@@
  • We propose a novel unified monitoring and forecasting framework using video diffusion models, and it inherently supports uncertainty analysis.
@@
The remainder of the paper is organized as follows. In section 2, we introduce the video diffusion models and the proposed framework for forecasting and monitoring based on video diffusion models with reconstruction guidance. Then, we share our numerical simulation for the data generation. In section 3, we present numerical tests of the proposed method on compass models, including the results of unconditional generation, monitoring, forecasting, and uncertainty analysis. Finally, we discuss and summarize the work, and present insights in sections 4 and 5.
@@ 2.2 Video Diffusion models
Denoising diffusion probabilistic models (DDPMs) are a family of generative models that try to represent a data distribution which enables the generation of samples from the data distribution. They have shown a state-of-the-art image and video synthesis performance among other generative models, such as variational autoencoders [Kingma and Welling, 2013], generative adversarial networks [Goodfellow et al., 2014], normalizing flows [Rezende and Mohamed, 2015]. The general idea behind DDPMs is to train a neural network to learn to transform a prescribed distribution (often a Gaussian distribution) to an approximation of the target data distribution. The pipeline of the diffusion model includes two processes: training (a.k. forward process from $\mathbf{x}_{0}$ to $\mathbf{x}_{T-1}$) and sampling (a.k. reverse process from $\mathbf{x}_{T-1}$ to $\mathbf{x}_{0}$), shown in Figure 1. The subscript $T$ of the $\mathbf{x}_{T}$ here denotes the time step in the diffusion process. Our target is to fit the data distribution of spatial-temporal evolution for multiple variables, which are 3D volumes. We process it as videos and thus use video diffusion models [Ho et al., 2022].
@@
Figure 1: Workflow of the diffusion process and its reverse process, and the reconstruction guided generation for forecasting and inversion. The yellow box denotes the scaled learned score function. “History” denotes the given conditions $\mathbf{x}^{0:t}$ for the forecasting tasks to predict $\mathbf{x}^{t:T}$, where superscript stands time-lapse frames. “Observation” denotes partial physical variables of $\mathbf{x}^{0:T}$, e.g., RTM images, for the inversion of other variables.
@@ 2.5 CO2 related Multiphysics datasets generation
Fluid flow simulation. Before we can simulate fluid flow, we need to define the permeability $K$ of the subsurface model. Here, as suggested by Yin et al. [2023b], Louboutin et al. [2023b], we convert the wave speed $v$ of the compass model to log-permeability values to make up alternating high- and low-permeability layers in the reservoir with a seal on top. The conversion follows a non-linear relationship, where the permeability $K$ is defined as follows: $K=3000e^{(v-4)}$ for $v\geq 4$; $K=0.01e^{(25.223(v-3.5))}$ for $43.5\leq v<$ 4; and $K=0.01e^{(v-3.5)}$ for $v<3.5$. This permeability is generated once based on velocity and kept constant regardless of the velocity changes due to CO2 injection. The porosity $\phi$ is also set to a constant value of 0.25. Then, we place one injection well in the deeper part of the target domain and consider a constant injection rate for the experiment here. We use a two-phase simulation to obtain the CO2 saturation model in Figure 2 with softwares Jutul.jl [Møyner et al., 2023] and JutulDarcyRules.jl [Yin et al., 2023a].
Rock physics modeling. Since we use seismic methods to monitor the subsurface CO2 changes, we need to obtain the velocity changes resulting from injection and the changes in saturation. We use the patchy saturation model [Louboutin et al., 2023b] to convert the saturation model to velocity and density models, in which an increase in CO2 saturation will result in a decrease in velocity and density. We ignore the pressure effect here and assume the shear wave velocity is fixed as $\frac{v}{\sqrt{3}}$, and the non-linear conversion follows
$\begin{array}[]{ll}\hat{\rho}&=\rho-0.276\phi S\\ \hat{v}&=\sqrt{{\frac{1}{{\hat{\rho}}}\left[(1-S)\left(\rho v^{2}\right)^{-1}+S\left(\frac{B_{min}}{(\par\frac{\rho v^{2}}{1.8B_{min}-\rho v^{2}}-\par\frac{2.735}{\phi\left(B_{min}-2.735\right)}+\par\frac{0.125}{\phi\left(B_{min}-0.125\right)}\par)^{-1}+1}+\frac{4}{9}\rho v^{2}\right)^{-1}\right]^{-1}}}\end{array}$ (7)
[…]
Seismic modeling and reverse time imaging. After obtaining the seismic properties changes represented by the velocity and density, we simulated the seismic recordings based on the impedance model calculated from the velocity and density, and through using born modeling. We then applied RTM to obtain seismic imaging results [Yin et al., 2023b] by the software JUDI.jl [Louboutin et al., 2023a]. The details on the source and receiver locations will be shared in the examples section. Here, we use RTM images as the seismic response for convenience in testing our concept. As the seismic responses to the velocity and density changes are quite small compared to the baseline monitoring, we do not use the RTM images directly. Instead, we use the residuals between the current RTM image and the baseline RTM image as the seismic observation. The Born modeling here will secure for us clean images as if processing and demultiple were applied to conventionally observed data.
@@ 2.6 The backbone network and training configuration
The backbone of the video diffusion model can have any architecture, e.g., 3D-Unet based [Cicek et al., 2016] or transformer based [Peebles and Xie, 2022]. In this paper, since we do not use a variational autoencoder to compress the videos to the latent space, using a transformer would be costly due to the quadratic computational complexity in the attention mechanism of the transformer. Thus, we use a 3D-Unet, which is factorized over space and time. Compared to the 2D Unet often used for image processing, whose convolutional layer is often with the kernel size of 3$\times$3, we use 1$\times$3$\times$3 convolutional kernel for the video. The dimensions of the kernel correspond to the time frames, height, and width of the multiphysics evolution. The attention in spatial attention block used in the 3D-Unet applies to the space dimension only, in which the time frame axis is regarded as the batch dimension, combined with a following temporal attention block. To distinguish the order of the frames (the time order of the multiphysics evolution), the relative position embeddings [Shaw et al., 2018] are used in each temporal attention block. This paper’s backbone network of the video diffusion model is designed to consider five channels, including permeability, CO2 saturation, velocity, density, and RTM images.
@@ 3 Numerical Experiments
In our experiments, we employed the 3D Compass velocity model, a synthetic model that closely mimics the geological characteristics found in the southeastern region of the North Sea. This model serves as an ideal testbed due to its realistic representation of the subsurface geological features. Since we focus more on the demonstration of the framework, we extract 514 two-dimensional slices from the central deep part of the three-dimensional compass velocity model, having dimensions of $64\times 64$, with an interval of 25 $m$ in the horizontal direction and 12 $m$ in the vertical direction. In this paper, we do not consider the leakage scenario and assume the permeability is constant over time, though this scenario can also be incorporated if needed. For the fluid flow simulation, we control the CO2 injection rate to achieve a 10% storage capacity over an 18-year period. The constant CO2 injection happens at the same fixed well location for all cases (often known in field experiments). From this simulation, we sampled six frames starting from the third year with a time-lapse interval of three years. These frames were crucial in analyzing the temporal evolution of the injected CO2. Subsequently, we applied rock physics modeling to calculate the corresponding changes in velocities and densities due to CO2 injection for each CO2 saturation at different times. We collected the simulated seismic data for each velocity and density model by uniformly placing six sources on the surface with a 24 Hz Ricker wavelet and 200 receivers on the sea bottom. The recordings lasted for 1 $s$ with a sampling rate of 4 $ms$. Then, we employ RTM to compute the time-lapse seismic image changes by subtracting them from the baseline image. In the video diffusion training phase, which is conducted on two Nvidia A100 GPUs for 32.8 hours, we used an Adam optimizer with a learning rate of 5e-5, a batch size of 128, and a maximum of 50,000 iterations until the neural network converged, as evidenced by the training loss curve, shown in Figure 3.
@@ 3.2 Forecasting
As mentioned in Section 2, this well-trained unconditional video diffusion model can be used for forecasting and inversion. We first test it for the forecasting task to evaluate the ability to predict future CO2 related multiphysics evolution based on the history frames as that can help us evaluate the capacity for CO2 storage for decision-making optimization. With a reconstruction scale $w_{r}$ of $10^{6}$, the conditional sampling results are shown in Figure 5. Given the observations and records of the subsurface states in years 3, 6, 9, we forecast the future states for years 12, 15, and 18. They are consistent and reasonable in their evolution and generally maintain the coupling relationship between the five physical parameters. To quantitate the forecasting errors for each variable, we compare them with the reference numerical simulations, shown in Figure 6. The errors are small and are focused mainly in the region of the CO2 plume. This generally validates our model’s accuracy and highlights its potential as a powerful tool for strategic planning in CO2 storage projects. We test the forecasting performance on five distinct examples, which are not used during the training. The average relative L2 norm error, which is calculated by dividing the L2 norm of the error by the L2 norm of the ground truth and the lower value denotes a better performance, is 0.024, and the average structural similarity index measure (SSIM, the higher value denotes a better performance) [Zeng and Wang, 2012] is 0.889.
@@
Figure 5: Forecasting generation (b) given three history frames (a), which are denoted by the red box. The solid color subfigures in (a) denote the unknown future states, which are the target of the forecasting generation. The rows represent properties described in Figure 4.
@@
Figure 7: Inversion results (b) given three history frames and observed-data-based RTM images (a), which are denoted by the red box. The solid color subfigures in (a) denote the unknown partial physical variables, which are the target of the inversion. The rows represent properties described in Figure 4